OpenStudy (anonymous):

Factor completely x^2-4x-21

OpenStudy (anonymous):

@Vlery

OpenStudy (anonymous):

What times what gives you 21 and also can give you the sum of -4? 7 * 3 = 21 -7 + 3 = -4 so our factors are 3 and -7 so it is factored to (x -7) (x + 3)